Solve. 6x + 4(3x – 2) = 6x + 2
Answer by tutorcecilia(2152)

6x + 4(3x – 2) = 6x + 2
6x + 12x – 8 = 6x + 2
18x – 8 = 6x + 2
18x -6x – 8 = 6x -6x+ 2
12x – 8 = 2
12x – 8 + 8 = 2 + 8
12x = 10
12x/12 = 10/12
x = 5/6 [Final answer]
.
.
6x + 4(3x – 2) = 6x + 2 [Check by plugging in (x= 5/6) back into the original equation]
6(5/6) + 4[(3)(5/6) – 2)] = 6(5/6) + 2 [Simplify]
5 – 4[(15/6) - 2] = 5 + 2
5 + (60/6) – 8 = 7
5 + 10 – 8 = 7
7=7 [checks out]